OpenStudy (anonymous):

can any one show me how to do 33% off 48.95 ??:)

OpenStudy (anonymous):

48.95 x 0.33 = 16.1535

OpenStudy (anonymous):

Because technically 33% is the same thing as 0.33.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

33% off would imply 33% removed from the original amount, not 33% of the original amount. That calculation would be (1-0.33)*48.95=0.67*48.95=32.7965

OpenStudy (anonymous):

True true. Sorry!